OpenStudy (anonymous):

Can you check my answer? Nakim simplified 3√2x + x√8x - 5√18x and got the answer -10√2x Part 1:Using complete sentences, explain what Nakim did wrong. Part 2: Show all your work to simplify the expression. Show all your work to simplify the expression. My response was: 3√2x cannot be simplified any further. x√8x = x√2*2*2x = 2x√2x -5√18x = -5√3*3*2x = -5*3√x = -15√x Combine like terms 3√2x + 2x√2x = 5√2x - 15√x = -15√x So my answer is 5√2x - 15√x

hartnn (hartnn):

like terms are 3√2x and - 15√x

hartnn (hartnn):

2x√2x remains as it is

hartnn (hartnn):

also, -5√18x is not simplified properly.

hartnn (hartnn):

-5√18x = -5√3*3*2x correct till here.

hartnn (hartnn):

-5√18x = -5√3*3*2x = -5*3 *√2x right ?

OpenStudy (phi):

you said My response was: 3√2x cannot be simplified any further. <--- true x√8x = x√2*2*2x = 2x√2x <------------true -5√18x = -5√3*3*2x = -5*3√x = -15√x <--- you lost a 2 that should be -15 sqrt(2x)

hartnn (hartnn):

@haileemackk you there ?

OpenStudy (anonymous):

Sorry, I had to leave after I posted this and I left my computer on. Even when I combined 3√2x + 2x√2x - 15√2x = -10√2x I still got -10√2x But if 2x√2x is on it's own, does that mean the answer is 3√2x - 15√2x = -12√2x + 2x√2x?

OpenStudy (phi):

maybe you see what is going on this way. factor out \( \sqrt{2x} \) from each term to get \[ (3 -15 + 2x) \sqrt{2x} \] or \[ (2x -12) \sqrt{2x} \] you cannot simplify 2x+12. we could factor out a 2 and write it as \[ (x-12) 2 \sqrt{2x} \] or we can write it as you did \[ 2x \sqrt{2x} -12 \sqrt{2x} \] any of those are ok, and any might show up as an answer in a multiple choice.

OpenStudy (phi):

** \[ (x-6) 2 \sqrt{2x} \] typo on one of the versions

OpenStudy (anonymous):

When you factored it out, I understood much better! Thank you so much!